# Reversal of Thought: Enhancing Large Language Models with Preference-Guided Reverse Reasoning Warm-up

### v2

@@ Abstract
Jiahao Yuan1, Dehui Du1, Hao Zhang1, Zixiang Di1, Usman Naseem2 1East China Normal University 2Macquarie University 51275900024@stu.ecnu.edu.cn, dhdu@sei.ecnu.edu.cn, {71275902039,51265901113}@stu.ecnu.edu.cn, usman.naseem@mq.edu.au
@@ 1 Introduction
Recent studies have advanced CoT to guide LLMs, mainly through either multi-step prompting such as introducing planning-and-solve Plaat et al. (2024); Yang et al. (2024a), self-consistency Narang et al. ; Wang et al. (2024) and recursive reasoning process Lee and Kim (2023); Yu et al. (2024) through Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024), or multi-role Zhang et al. ; Suzgun and Kalai (2024) to enhance logical capabilities and mitigate hallucination, yet this has stealthily increased inference cost due to the multi-step inference. Buffer-of-Thought (BoT) Yang et al. (2024a) attempts to reduce thinking steps by leveraging Retrieval-Augmented Generation (RAG) to retrieve gold thought templates from the buffer. However, it sacrifices flexibility due to the initialization of pre-set manual thought templates. Therefore, achieving accurate reasoning in LLMs while minimizing resource consumption remains a significant challenge.
@@ 2.1 Chain-of-Thought (CoT) Prompting
Chain-of-Thought (CoT) prompting Wei et al. (2022) has been proven to be a promising approach that incorporates an intermediate logic chain to enhance LLMs’ logic. Recent studies primarily aimed at improving logical accuracy by external validation mechanisms like symbolic reasoning Cai et al. (2023); Pan et al. , stepwise verification including self-consistency Narang et al. ; Yu et al. (2024); Wang et al. (2024), self-refine Madaan et al. (2024), self-reflection Renze and Guven (2024) and more hierarchical information such as Least-to-Most Zhou et al. , Cumulative-Reasoning Zhang et al. and Multi-experts Suzgun and Kalai (2024) strategies, but faced challenges related to cumulative errors Bao et al. (2024) or poor flexibility Yang et al. (2024a). Additionally, numerous studies also proposed more standardized recursive or backtracking branch forms from the logical data structure, including Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024) and Buffer-of-Thought (BoT) Yang et al. (2024a). However, an efficient logical reasoning method that strikes a balance among reasoning accuracy, flexibility, and cost has yet to be discovered. Our method is activated through meta cognition Fleur et al. (2021) by introducing reverse reasoning to form effective LLMs-taste prompts within cognitive preference Uddin (2021) for plan-and-solve with logical pseudocode at least.
@@ 3.2 Reverse Reasoning with Meta-cognition
Following the principle of preference transitivity Liu et al. (2024b), we extend $P(R_{i+1}\succ R_{i})$ to $P(R_{i}\succ R_{j})$ to reduce computational cost from, thereby forming a preference matrix $P_{pre}\in\mathbb{R}^{warm\times warm}$, formally:
@@ 4.1 Datasets and Tasks
Following Yao et al. (2024); Suzgun and Kalai (2024); Yang et al. (2024a), we evaluate RoT across eight tasks (detailed in Appendix A.1): (1) Mathematical Reasoning:Game of 24 Yao et al. (2024); Xiang et al. (2025), Multilingual Grade School Math Shi et al. (2022); (2) Domain Knowledge: Python Puzzles Schuster et al. (2021), Geometric Shapes, Multi-Step Arithmetic Two, Word Sorting, and Checkmate-in-One Srivastava et al. (2023); (3) Creativity: Shakespearean Sonnet Writing Suzgun and Kalai (2024).
@@ 4.3 Experiment Setup
To ensure fair comparisons with previous methods, following Liu et al. (2024a); Yao et al. (2024); Zhang et al. ; Suzgun and Kalai (2024), we utilize two LLMs: GPT-3.5-turbo and GPT-4, as the foundational model via OpenAI API for our RoT experiments including both main experiments and ablation study. For the $warmup$ hyperparameter, we experimented with values of $\{1,3,5,10\}$, accessed in batches for reverse reasoning warm-up through OpenAI API. Our findings suggest that a value of 5 optimally balances between logical accuracy and cost-efficiency. For embedding model $\mathcal{M}_{emb}$, we utilize a huggingface model stella_en_1.5B_v5  [^2] https://huggingface.co/dunzhang/stella_en_1.5B_v5, License:mit, a high-performance model with the smallest parameter countamong the top three on the MTEB leaderboard Muennighoff et al. (2023), offline deployed on a single NVIDIA GeForce RTX 4090 GPU.
@@ 4.4 Evaluation Metrics
Following Suzgun and Kalai (2024); Yao et al. (2024), we introduce a LLM444we select the latest openai-o1 as judger, https://openai.com/index/introducing-openai-o1-preview/ to validate the final logical reasoning against gold results (Correct/Wrong). We then compute logical accuracy $Acc_{logic}$ for each logical task by tallying the number of correct responses.
@@ 5.1 Knowledge Boundary and Cognitive Preference Consistency
Knowledge Boundary.
Cognitive Preference Consistency.
5.2 Reasoning Accuracy and Efficiency
(1) RoT can activate LLMs’ reasoning accuracy.
Figure 3: Inference time comparison
@@ 6 Conclusion
In future work, we aim to extend RoT’s capabilities by incorporating In-Context Learning (ICL), which will allow for greater flexibility in adapting to varied contexts and improve its performance on more complex reasoning tasks. Furthermore, we believe that utilizing Reversal of Thought in teacher-student model distillation, which could further amplify the practical value of our approach.
@@ Appendix A Experimental Details
A.1 Datasets and Tasks
To comprehensively validate our Reverse of Thought (RoT), following ToT Yao et al. (2024), meta-prompting Suzgun and Kalai (2024) and BoT Yang et al. (2024a), we assess it with baselines across a broad range of eight tasks across five logical benchmarks that encompass mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity:(1) Game of 24 Yao et al. (2024); Xiang et al. (2025) challenges LLMs to create a mathematical expression utilizing each of four given numbers exactly once to achieve 24666https://www.4nums.com/game/ . (2) BIG-Bench Suzgun et al. (2023); Srivastava et al. (2023) involves Geometric Shapes, Multi-Step Arithmetic Two and Word Sorting and Checkmate-in-One from BIG-Bench suite  [^7] https://huggingface.co/datasets/google/bigbench Srivastava et al. (2023); (3) Python Puzzles Schuster et al. (2021) comprises a collection of challenging programming puzzles crafted in Python, covering various difficulty levels [^8] https://github.com/microsoft/PythonProgrammingPuzzles;(4) Multilingual Grade School Math (MGSM) Shi et al. (2022) is a multilingual adaptation of the GSM8K dataset Cobbe et al. (2021), featuring translations of a subset of examples into ten diverse languages [^9] https://github.com/google-research/url-nlp/tree/main/mgsm .(5) Shakespearean Sonnet Writing Suzgun and Kalai (2024) require LLMs to compose with the rhyme scheme "ABAB CDCD EFEF GG" while incorporating three  [^10] https://huggingface.co/datasets/turingmachine/meta-prompting or five  [^11] https://github.com/iljones00/Shakespearean-Sonnets-GPT specified words verbatim.
A.2 Prompt for Reverse Reasoning
@@ 3. Formulate Logical Pseudocode:
A.3 Prompt for CPM (Known/Unknown)
@@ 7. Fourth train arrival time = Add(Third train departure time, $\varepsilon$) = $125+20=145$ minutes
Appendix C Instructions for Human Evaluation.
To evaluate the cognitive preference consistency $Con_{cog}$ between LLMs-taste prompt for task cognitions across different tasks, annotators adhere to the following guidelines:
@@
Responses are evaluated on a scale from 0 to 5. A score of 4-5 signifies exemplary style consistency, while 3-4 indicates strong consistency with minor variations. Scores of 2-3 represent adequate style but with noticeable inconsistencies, and 1-2 denote inconsistent style across tasks. A score of 0-1 reflects complete inconsistency.

### v3

@@ Abstract
Jiahao Yuan1thanks: 51275900024@stu.ecnu.edu.cn, Dehui Du1thanks: Corresponding Author: dhdu@sei.ecnu.edu.cn, Hao Zhang1, Zixiang Di1, Usman Naseem2 1East China Normal University 2Macquarie University
@@ 1 Introduction
Recent studies have advanced CoT to guide LLMs, mainly through either multi-step prompting such as introducing planning-and-solve Plaat et al. (2024); Yang et al. (2024a), self-consistency Narang et al. ; Wang et al. (2024a) and recursive reasoning process Lee and Kim (2023); Yu et al. (2024) through Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024), or multi-role Zhang et al. ; Suzgun and Kalai (2024) to enhance logical capabilities and mitigate hallucination, yet this has stealthily increased inference cost due to the multi-step inference. Buffer-of-Thought (BoT) Yang et al. (2024a) attempts to reduce thinking steps by leveraging Retrieval-Augmented Generation (RAG) to retrieve gold thought templates from the buffer. However, it sacrifices flexibility due to the initialization of pre-set manual thought templates. Therefore, achieving accurate reasoning in LLMs while minimizing resource consumption remains a significant challenge.
@@ 2.1 Chain-of-Thought (CoT) Prompting
Chain-of-Thought (CoT) prompting Wei et al. (2022) has been proven to be a promising approach that incorporates an intermediate logic chain to enhance LLMs’ logic. Recent studies primarily aimed at improving logical accuracy by external validation mechanisms Bi et al. (2025) like symbolic reasoning Cai et al. (2023); Pan et al. , stepwise verification including self-consistency Narang et al. ; Yu et al. (2024); Wang et al. (2024a), self-refine Madaan et al. (2024), self-reflection Renze and Guven (2024) and more hierarchical information such as Least-to-Most Zhou et al. , Cumulative-Reasoning Zhang et al. and Multi-experts Suzgun and Kalai (2024) strategies, but faced challenges related to cumulative errors Bao et al. (2024) or poor flexibility Yang et al. (2024a). Additionally, numerous studies also proposed more standardized recursive or backtracking branch forms from the logical data structure, including Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024) and Buffer-of-Thought (BoT) Yang et al. (2024a). However, an efficient logical reasoning method that strikes a balance among reasoning accuracy, flexibility, and cost has yet to be discovered. Our method is activated through meta cognition Fleur et al. (2021) by introducing reverse reasoning to form effective LLMs-taste prompts within cognitive preference Uddin (2021) for plan-and-solve with logical pseudocode at least.
@@ 3.2 Reverse Reasoning with Meta-cognition
Following the principle of preference transitivity Liu et al. (2024), we extend $P(R_{i+1}\succ R_{i})$ to $P(R_{i}\succ R_{j})$ to reduce computational cost from, thereby forming a preference matrix $P_{pre}\in\mathbb{R}^{warm\times warm}$, formally:
@@ 4.1 Datasets and Tasks
To comprehensively validate our Reverse of Thought (RoT), following ToT Yao et al. (2024), meta-prompting Suzgun and Kalai (2024) and BoT Yang et al. (2024a), we assess it with baselines across a broad range of eight tasks across five logical benchmarks that encompass mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity: (1) Game of 24 Yao et al. (2024); Xiang et al. (2025) challenges LLMs to create a mathematical expression utilizing each of four given numbers exactly once to achieve 24  [^2] https://www.4nums.com/game/ . We use 1000 samples drawn from a mix of both datasets for robust large-scale evaluation. (2) BIG-Bench Suzgun et al. (2023); Srivastava et al. (2023) involves Geometric Shapes, Multi-Step Arithmetic Two and Word Sorting and Checkmate-in-One from BIG-Bench suite  [^3] https://huggingface.co/datasets/google/bigbench Srivastava et al. (2023). (3) Python Puzzles Schuster et al. (2021) comprises a collection of challenging programming puzzles crafted in Python, covering various difficulty levels  [^4] https://github.com/microsoft/PythonProgrammingPuzzles. (4) Multilingual Grade School Math (MGSM) Shi et al. (2022) is a multilingual adaptation of the GSM8K dataset Cobbe et al. (2021), featuring translations of a subset of examples into ten diverse languages [^5] https://github.com/google-research/url-nlp/tree/main/mgsm. (5) Shakespearean Sonnet Writing Suzgun and Kalai (2024) require LLMs to compose with the rhyme scheme "ABAB CDCD EFEF GG" while incorporating three  [^6] https://huggingface.co/datasets/turingmachine/meta-prompting or five  [^7] https://github.com/iljones00/Shakespearean-Sonnets-GPT specified words verbatim.
@@ 4.3 Experiment Setup
To ensure fair comparisons with previous methods, following Yao et al. (2024); Zhang et al. ; Suzgun and Kalai (2024), we utilize two LLMs: GPT-3.5-turbo and GPT-4, as the foundational model via OpenAI API for our RoT experiments including both main experiments and ablation study. For candidate generation, we use a temperature of 0.7, encouraging diversity in proposed solution. For instantiating reasoning steps, we adopt a lower temperature of 0.1, promoting more deterministic and stable logical progression. For the $warmup$ hyperparameter, we experimented with values of $\{1,3,5,10\}$, accessed in batches for reverse reasoning warm-up through OpenAI API. Our findings suggest that a value of 5 optimally balances between logical accuracy and cost-efficiency. For embedding model $\mathcal{M}_{emb}$, we utilize a huggingface model stella_en_1.5B_v5  [^8] https://huggingface.co/dunzhang/stella_en_1.5B_v5, License:mit, a high-performance model with the smallest parameter countamong the top three on the MTEB leaderboard Muennighoff et al. (2023), offline deployed on a single NVIDIA GeForce RTX 4090 GPU.
Knowledge Boundary and Cognitive Preference Consistency.
@@ 4.4 Evaluation Metrics
Following Suzgun and Kalai (2024); Yao et al. (2024), we introduce a LLM101010we select the latest openai-o1 as judger during the development of our work, https://openai.com/index/introducing-openai-o1-preview/ to validate the final logical reasoning against gold results (Correct/Wrong). We then compute logical accuracy $Acc_{logic}$ for each logical task by tallying the number of correct responses.
@@ 5.1 Knowledge Boundary and Cognitive Preference Consistency
5.2 Reasoning Accuracy and Efficiency
(1) RoT can activate LLMs’ reasoning accuracy.
Figure 3: Inference time comparison, measured as the average duration from inference start to evaluation end, including all steps.
@@ 6 Conclusion
In future work, we aim to extend RoT’s capabilities by incorporating In-Context Learning (ICL) or, alternatively, adopting Auto-prompt based demonstration selection strategies Zhang et al. (2023); Jin and Lu (2024), which will allow for greater flexibility in adapting to varied contexts and improve its performance on more complex reasoning tasks. Furthermore, we believe that utilizing Reversal of Thought in teacher-student model distillation, which could further amplify the practical value of our approach.
Finally, we acknowledge the challenges posed by the inherent instability of large language models. Future research may benefit from investigating robust candidate solutions—such as hybrid scoring mechanisms Wang et al. (2024b); Minh et al. (2024); Wang et al. (2024c)—that can be integrated with RoT to further enhance its reliability and overall performance through joint optimization.
@@ Appendix A Experimental Details
A.1 Prompt for Reverse Reasoning
@@ 3. Formulate Logical Pseudocode:
A.2 Prompt for CPM (Known/Unknown)
@@ B.1.2 Instantiation Case Study for Game of 24 Using $P_{final}$
Figure 7: Instantiation Case Study for Game of 24 Using $P_{final}$
@@ 7. Fourth train arrival time = Add(Third train departure time, $\varepsilon$) = $125+20=145$ minutes
Figure 9: Instantiation Case Study for MGSM Using $P_{final}$
Appendix C Instructions for Human Evaluation
To evaluate the cognitive preference consistency $Con_{cog}$ between LLMs-taste prompt for task cognitions across different tasks, three professional annotators adhere to the following guidelines:
@@ Appendix C Instructions for Human Evaluation
Responses are evaluated on a scale from 0 to 5. A score of 4-5 signifies exemplary style consistency, while 3-4 indicates strong consistency with minor variations. Scores of 2-3 represent adequate style but with noticeable inconsistencies, and 1-2 denote inconsistent style across tasks. A score of 0-1 reflects complete inconsistency. Prior to the formal evaluation, several pre-scored sample responses covering every score band will be provided. These examples serve to calibrate annotators’ judgments and to align their understanding of the rubric.